5. Juli 2004. U. Schoenwaelder; http://www.math.rwth-aachen.de/∼Ulrich.Schoenwaelder HB = Hochschulbibl. RWTH, HBZ = http://www.hbz-nrw.de/ (HBZ-CD-ROM Online), MB = Mathematikbibl., DB = Didaktikbibl. (Winter), EB = Erziehungswiss. Bibl., FH = Bibl. Fachhochschule Aachen, FL = Fernleihe, IB = Informatikbibliothek, IB Nr. = Institutsbibliothek Nr., LB = HB– Lehrbuchsammlung, LS = HB–Lesesaal

Literatur zu p–adischen Zahlen

[1] W. A. Adams and L. J. Goldstein. Introduction to Number Theory. Prentice–Hall, Inc., 1976. MB: 9094. Quadratic fields and forms. S. 66–85: Solving Polynomial Congruences (p-adic). Chapters: 1 Introduction; 2 Divisibility and Primes; 3 Congurences; 4 The Law of Quadratic Reciprocity; 5 Arithmetic Functions (5.3 The Möbius Inversion Formula); 6 A Few Diophantine Equations; 7 The Gaussian Integers; 8 Arithmetic in Quadratic Fields; 9 Factorization Theory in Quadratic Fields; 10 Applications of the Factorization Theory to Diophantine Equations; 11 The Representation of Integers by Binary Quadratic Forms.

[2] G. Bachman. Introduction to p–adic Numbers and Valuation Theory. Academic Paperbacks. Acad. Press, 1964. MB: 3508.

[3] S. I. Borewicz and I. R. Shafarevich. Zahlentheorie. Birkhäuser Verlag, 1966. MB: 3600, HB: 32 Za 5775. Kap. I §3:p– adische Zahlen; §4: Axiomatische Charakterisierung des Körpers der p–adischen Zahlen.

[4] S. Bosch, U. Güntzer, and R. Remmert. Non–Archimedean Analysis. A Systematic Approach to Rigid Analytic Geometry. Grundlehren der mathematischen Wissenschasften 261. Springer–Verlag, 1984. HB: 261^{1} Bb 1285, MB: 12241.

[5] J. W. S. Cassels. Local Fields. London Mathematical Society Student Texts 3. Cambridge University Press, 1986. MB: 13222.

[6] A. A. Cuoco. Visualizing the p–adic integers. Amer. Math. Monthly, 98(4):355–364, 1991. MB: Z 42.

[7] H.-D. Ebbinghaus, H. Hermes, F. Hirzebruch, M. Koecher, K. Mainzer, J. Neukirch, A. Prestel, and R. Remmert. Zahlen. Springer–Verlag, 1983, ^{2}1988. MB: 12139, 12165, 14614. Es gibt eine englische Übersetzung. 2. Aufl.: J. Neukirch, p– adische Zahlen.

[8] H. M. Edgar. A First Course in Number Theory. Wadsworth, 1988. p–adic Newton’s method for successive approximations. Tel. Rev.: Amer. Math. Monthly 99:8(1992), 805.

[9] G. Frey. Elementare Zahlentheorie. Vieweg–Studium 56, Grundkurs Mathematik. Vieweg, 1984. MB: 13930. Kap. III, §4: p–adische Zahlen.

[10] A. Fröhlich and M. J. Taylor. Algebraic Number Theory. Cambridge Studies in Advanced Mathematics 27. Cambridge University Press, 1991. MB: 16171. I §2: Valuations and absolute values; II §3: Completions.

[11] F. Gouvêa. p–adic Numbers. Universitext. Springer–Verlag, 1993. MB: 17152.

[12] N. Koblitz. p–adic Numbers, p–adic Analysis, and Zeta–Functions. GTM 58. Springer–Verlag, 1977. MB: 9544. HB: 58 Bb 1244.

[13] H. Koch. Algebraic number fields. In A. N. Parshin and I. R. Shafarevich, editors, Number Theory II, Encyclopedia of Mathematical Sciences 62. Springer–Verlag, 1992. MB: 15893 b. Übersicht. Umfangreiches Literaturverzeichnis. Ch. I §4: Valuations.

[14] S. Lang. Undergraduate Algebra. UTM. Springer–Verlag, 1987,^{2}1990. MB: 13625. S. 288 (Exercises): p–adic integers as projective limit; S. 289–290: completion of p–adic integers.

[15] A. Leutbecher. Zahlentheorie. Eine Einführung in die Algebra. Grundwissen Mathematik, Springer–Lehrbuch. Springer– Verlag, 1996. MB: 17917. Ch. 9: Die lokalen Körper über Q. Ch. 20: Die Bewertungen der Zahlkörper.

[16] W. J. LeVeque. Fundamentals of Number Theory. Addison–Wesley, 1977. MB: 9885. Ch. 3.5: The p–adic fields.

[17] G. Meyerson and A. J. van der Poorten. Some problems concerning recurrence sequences. Amer. Math. Monthly, 102(8):698–705, 1995. §4: An introduction to p–adic analysis.

[18] Deborah Nolan, editor. Women in Mathematics. Scaling the Hights, MAA Notes. MAA, 1997. Ein Topic: p–adic numbers. For Undergraduates.

[19] V. P. Platonov and V. I. Yanchevskii. Finite-dimensional division algebras. In A. I. Kostrikin and I. R. Shafarevich, editors, Algebra IX, Encyclop. Math.Sci. 77, pages 121–233. Springer–Verlag, 1996. MB: 15143i. S. 130–133: Ch. 1.5. Fields of formal Laurent series and p–adic numbers.

[20] L. Rédei. Endliche p–Gruppen. Akadémiai Kiadó, Budapest, 1989. MB: 14755. Kap. I: Der p–adische Zahlkörper und die Grundlagen der p-adischen Analysis.

[21] Alain M. Robert. A Course in p-Adic Analysis. Graduate Texts in Mathematics 198. New York: Springer, 2000. ISBN 0-387-98669-3. MB: 18940. HBZ. Review: Bulletin AMS 39:1 (2002), 113-119.

[22] Joseph J. Roseman. Second order iterations. The College Mathematics Journal, 30(5):393–396, 1999. Konvergenzordnung. Illustrates the concept of iteration through an example. Example: iterative division (for numbers in a mixed decimal-2adic form). We want a formula of the form x_{n+1} = g(x_{n}).

[23] M. I. Rosen. Niels Hendrik Abel and equations of the fifth degree. Amer. Math. Monthly, 102(6):495–505, 1995. MB: Z 42. S. 505: Anwendung des Weak Aprroximation Theorem und Hensels Lemma zum Nachweis der Existenz eines Q-Polynoms vom Primzahlgrad, dessen Nullstellen alle reell sind, das aber nicht durch Radikale lösbar ist.

[24] J.-P. Serre. Cours d’Arithmétique. Presses Universitaires de France, Paris, 1970. Vgl. [25].

[25] J.-P. Serre. A Course in Arithmetic. GTM 7. Springer–Verlag, 1973. HB: 7 Bb 1244. Vgl. [24].

[26] G. Stroth. Algebra. Einführung in die Galoistheorie. de Gruyter Lehrbuch. de Gruyter, 1998. MB: 18819. §14: p–adische Zahlen (mit Hensellemma).

[27] Hermann Weyl. Algebraische Zahlentheorie. Hochschultaschenbücher 136/136a. Mannheim: BI, 1966. MB: 4420. Kap. III: Lokale p-adische Untersuchungen (Kummer, Hensel). Englisches Original: Algebraic Theory of Numbers, 1940, MB: 486.